OpenStudy (anonymous):

Find the equation of the axis of symmetry and the coordinates of the vertex of the graph of the following function y=-3x^2+4x-5

OpenStudy (welshfella):

you need to convert this to the form a(x - b)^2 + c where the axis of symmetry is x = b

OpenStudy (welshfella):

-3x^2 + 4x - 5 divide through by -3:- = -3(x^2 - (4/3) x +5/3) = -3((x - (2/3 )^2 - 16/9 + 5/3) can you see what the axis of symmetry is ? - compare it with the general form in my first post

OpenStudy (welshfella):

look where b is and what is in the same place in the last line of my working

OpenStudy (welshfella):

going from x^2 - (4/3) x to (x - (2/3) x )^2 -16/9 is called completing the square

OpenStudy (anonymous):

so the axis of symmetry would be 2/3

OpenStudy (welshfella):

x = 2/3 yes